OpenStudy (anonymous):

The length of the diagonal of a square is 8 inches. What is the area of the square?

OpenStudy (maheshmeghwal9):

If the length of a side of a squre is 'a' then diagonal's length=\[a \sqrt{2}.\] so side= diagonal/sqrt{2} So area = (side)^2= (diagonal/sqrt{2})^2.

OpenStudy (maheshmeghwal9):

area=\[(\frac{diagonal}{\sqrt{2}})^2.\]

OpenStudy (anonymous):

so the answer is 32 ?

OpenStudy (maheshmeghwal9):

yeah well done:)